# Emergence of Chimeras States in One-dimensional Ising model with Long-Range Diffusion

### v1

@@ 1 Introduction
Our study shows that a simple Ising chain with non-local diffusion exhibits a rich phenomenology, including chimera states. In this scenario, we also show that these states constitute the equilibrium state satisfying detailed balance, for certain regions of the parameter space. The present work could be extended to some non-equilibrium situations, including, e.g., competing reaction and diffusion dynamics, a scenario more commonly observed in many social, ecological and biological systems. In this last case, a prominent example is the activity of the mammalian brain, where neural excitability is driven by chemical synapses (reaction) and voltage propagation occurs via electrical synapses (diffusion).
@@ 2.1 Non-local diffusion dynamics
$\mathcal{H}(\boldsymbol{\sigma})=-\frac{1}{2}\sum_{i,j}J_{ij}\sigma_{i}\sigma_{j}$ (2.1)
where $J_{ij}=J\times[0<|i-j|\leq R]$ (with $[P]$ indicating the Iverson bracket) indicates constant ferromagnetic coupling ($J>0$) within the diffusion range $R$. Note that $J$ is given in units of $k_{B}T.$ An illustration of the system is depicted in Fig. 1.
@@
Due to the presence of the thermal bath, the system evolves according to stochastic dynamics governed by a master equation with transition matrix $W(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma})$. Thus, letting $P(\boldsymbol{\sigma},t)$ be the probability of finding the system in state $\boldsymbol{\sigma}$ at time $t$, the evolution of the system in a characteristic time $\tau$ can be expressed as
$\tau\frac{dP(\boldsymbol{\sigma},t)}{dt}=\sum\limits_{\boldsymbol{\sigma}^{\prime}}[P(\boldsymbol{\sigma}^{\prime},t)W(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma})-P(\boldsymbol{\sigma},t)W(\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime})]$ (2.2)
$W(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma})=\sum\limits_{\begin{subarray}{c}i,j\\ 0<|i-j|\leq R\end{subarray}}\delta_{\sigma_{1},\sigma^{\prime}_{1}}...\delta_{\sigma_{i}\sigma^{\prime}_{j}}...\delta_{\sigma_{j}\sigma^{\prime}_{i}}...\delta_{\sigma_{N},\sigma^{\prime}_{N}}w^{ij}(\boldsymbol{\sigma})$ (2.3)
@@
$w^{ij}(\boldsymbol{\sigma})=\begin{cases}1\quad\mathrm{if}\;\Delta\mathcal{H}^{ij}\leq 0\\ e^{-\beta\Delta\mathcal{H}^{ij}}\quad\mathrm{if}\;\Delta\mathcal{H}^{ij}>0\ \end{cases}\quad\quad\Delta\mathcal{H}^{ij}=\mathcal{H}(\boldsymbol{\sigma}^{ij})-\mathcal{H}(\boldsymbol{\sigma})$ (2.4)
@@ 2.2 Simulations and numerical details
We now describe the Monte Carlo algorithm used to simulate our system for an extensive combination of model parameters. The Ising model itself has no intrinsic dynamic and it is when we couple the system to a heat bath at temperature $T$ that stochastic changes in the spin variables occurs so we have a stochastic kinetic version of Ising model driven by the master equation as those in Eq. 2.3. In a spin system of size $N$, each Monte Carlo Step (MCS) consists of $N$ spin-flip or spin exchange attempts. In general, we have performed simulations of $5\times 10^{4}$ MCS. Each simulation of our system operates according to the following algorithm:
@@
    The Metropolis rate $w^{ij}(\boldsymbol{\sigma})$ is then evaluated following the algorithm described in 2.4 and we compute $\Delta\mathcal{H}$ explicitly as follows:
    $\Delta\mathcal{H}^{ij}=\left(\sum\limits_{\begin{subarray}{c}\\ k\neq j\end{subarray}}J_{ik}\sigma_{k}-\sum\limits_{\begin{subarray}{c}k^{\prime}\neq j\end{subarray}}J_{k^{\prime}j}\sigma_{k^{\prime}}\right)\left(\sigma_{i}-\sigma_{j}\right)$ (2.5)
@@
Spin pairs in the ring are updated using a sequential updating. Although work regarding parallel updating (all spins updated in the ring at the same time step) in the Ising model has been conducted [42] and proven to obey equilibrium statistics, one has to be careful with its implementation so that it satisfies detailed balance [43].
@@ 3.1 Analytical results
$\Delta\mathcal{H}^{ij}=\left(\sum\limits_{\begin{subarray}{c}k=i-R\\ k\neq i,j\end{subarray}}^{i+R}\sigma_{k}-\sum\limits_{\begin{subarray}{c}k^{\prime}=j-R\\ k^{\prime}\neq i,j\end{subarray}}^{j+R}\sigma_{k^{\prime}}\right)\left(\sigma_{i}-\sigma_{j}\right)$ (3.1)
where we have just computed the terms in $\Delta\mathcal{H}$ summing over $R$ neighbors on each side and considered $J=1$ for simplicity. In other words, each candidate spin $\sigma_{i}$ “feels” the interaction of the sum of its $2R$ neighbors (omitting itself and $\sigma_{j}$), and each candidate $\sigma_{j}$ “feels” the interaction of the sum of its $2R$ neighbors (omitting itself and $\sigma_{i}$). Note that exchange energy $\Delta\mathcal{H}^{ij}$ is only non-zero if the exchanged spins $i$ and $j$ have different sign.
@@ 3.1.1 Emergence and stability of chimera states
$B_{n^{+}}(i_{1},\ldots,i_{n^{+}})=\left\{k\in\left\{1,\ldots,N\right\},\;k\in\bigcap_{m=1}^{n^{+}}B_{1}(i_{m}),\;|i_{1}-i_{n+}|\leq R\right\}$ (3.2)
@@
$p_{n^{+}}^{exchange}=\frac{2n^{+}}{N^{2}}|B_{n^{+}}(i_{1},...\,,i_{n^{+}})|\;\;\;\;\;\;\;\;\;\;\;\;p_{n^{+}}^{c}=1-p_{n^{+}}^{exchange}.$ (3.3)
@@
$W\bigl(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma}\bigr)\;=\;\frac{p_{n^{+}}^{exchange}}{n^{+}}\sum\limits_{k=1}^{n^{+}}\left[\sum\limits_{i_{k},j\in B_{n^{+}}(i_{1},...,i_{n^{+}})}Ex_{i_{k},j}(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma}\bigr)\right]+p_{n^{+}}^{c}\;\Delta(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma}\bigr)$ (3.4)
@@
$\varSigma^{n^{+}}:=\left\{\boldsymbol{\sigma}\in\Lambda\;/\;|i_{1}-i_{n^{+}}|_{p}\leq R,\;i_{1}<i_{2}<\ldots<i_{n^{+}}\right\}.$ (3.5)
@@
Also, note that the actual construction of $B_{n^{+}}(i_{1},...\,,i_{n^{+}})$ is not always possible, as can be see from definition 3.2. This follows from the fact that the interaction range $R$ limits the maximum number of positive spins that can simultaneously located within the range of another one. In fact, for configurations with $n^{+}>R+1\Rightarrow|i_{1}-i_{n^{+}}|_{p}>R\Rightarrow\;B_{n^{+}}=\emptyset$. In this situation, there is no rule to determine which spins will have zero energy transitions available. Actually, even for a configuration with $n^{+}=R+1$, regardless of whether $|i_{1}-i_{n^{+}}|_{p}\leq R$, the ball would still be the empty set. This is because any individual ball $B(i_{k})$ (from which we will construct $B_{n^{+}}$) is defined without the index $i_{k}$. This does not imply that exchanges are impossible, but that they are trivial as all spins are positive. We will try to discuss briefly what happens for $n^{+}>R+1$ in the next section.
@@ 3.1.2 Disappearance of chimera states and formation of attractors
Nonetheless, we have arrived to the energy minimum ($\boldsymbol{\sigma}\in\Sigma^{R}$), though any possible exchange would not change the configuration. Following equation 3.4:
@@
Let us now extend the number of positive spins considering now $n^{+}>R+1$. In this situation, $B_{n^{+}}=\emptyset$ as in the last case, but an important difference is that configurations in the energy minima will not be a configuration of type $\boldsymbol{\sigma}\in\Sigma^{n^{+}}$, since for $n^{+}>R+1$ one has $|i_{1}-i_{n^{+}}|_{p}>R$, so no configuration satisfies the condition to be in $\Sigma^{n^{+}}$, from definition 3.5. The system will anyway advance stochastically to an energy minimum, but this minimum will not be in any class of type $\Sigma^{X}$. However, we can give an argument of how these minima behave.
@@
To deduct the formation of these attractors, we have supposed that we have a configuration of type $\Pi$ from the start (see proposition 3.1.1). This will obviously not always be the case, and to characterize any kind of general behavior we have to employ a more general approach. In fact, we realized that the Hamiltonian (2.1) is a special case of a one‐dimensional Ising model with Kac‐type interactions [44, 45, 46]. For these, it is shown that regions of constant magnetization $m$ and m-m alternate throughout the system (what we have called attractors), and that they are of size $L_{attractor}\approx e^{cR}$ for some suitable constant $c$, which we will numerically calculate in the next section.
@@ 3.2 Numerical results for T=0
$m_{0}>\frac{2(R+1)}{N}-1=2r-1+{\cal{O}}\left(N^{-1}\right)=m_{c}$ (3.7)
Then, if we see that chimeras and attractors coexist in such conditions, this implies that that chimera state will be metastable, as chimeras will always be able to merge into attractors when 3.7 holds (see also next section).
@@ 3.2.1 Length of attractors
Fig. 3 plots $\ell_{dom}$ and $\tau_{ss}$ against $r$ for $m_{0}=0.4$. The figure illustrates that there are multiple $\ell_{dom}$ points for each value of $r$, which correspond to different-seeded runs. The scatter of low $\ell_{dom}$ points across nearly all $r$ values in Fig. 3 signals the intermittent presence of small, noise-driven domains, precisely the hallmark of chimera–attractor coexistence seen in Fig. 2. We fitted $\ell_{dom}$ versus $r$ to two candidate curves:
@@
Preliminary results in Fig. 4 show that the number of chimeras increase with $N$ and that through collisions they can form other chimeras or attractors. This last case occurs due to the fulfillment of condition 3.7 when $n^{+}$ is increased. So an interesting question to address is how the number of initials chimeras evolve in time until only attractors remain. We have seen that attractors start to emerge when $n^{+}\approx R+1.$ Note moreover that $\langle n^{+}\rangle\approx(1+m_{0})N/2\equiv q_{0}N$, Then, the condition for an attractor to emerge approximately is $q_{0}\approx(R+1)/N.$
@@
$N_{0}\approx\frac{q_{0}}{\gamma_{0}(R+1)}N\equiv\frac{\bar{\gamma}}{\gamma_{0}}N,$ (3.9)
@@
$L_{chimera}\approx R+1$ (3.10)
independently of $\gamma_{0}.$ This is consistent with definition in Eq. 3.6 where now and in general for any time one has $|B_{n^{+}_{c}}(i_{1},\ldots,i_{n^{+}_{c}})|=(1-\gamma(t))(R+1)$ and $n^{+}_{c}=\gamma(t)(R+1)$ We can conclude that initially
$N_{c}(0)=N_{0}\approx\frac{q_{0}}{\gamma_{0}L_{chimera}}N.$ (3.11)
@@
Under these assumptions, we have tested the Eq. 3.9 numerically in Fig. 5 (Main plot) for a relatively short observational time of $t_{obs}=1000$ MCS. The figure shows a clear scaling of the number of surviving chimeras at this time with the system size, while its inset indicates that when the observation time increases the number of existing chimeras decreases. Note, moreover, that the initial number of chimeras in the thermodynamic limit can already be established by Eq. 3.9 just taking $R=rN$ which yields $N_{0}\approx q_{0}/\gamma_{0}r$, expression that only depends on intensive quantities.
@@
$\tau_{c}\frac{dN_{c}(t)}{dt}=-2p_{0}f(\gamma(t))N_{c}(t)n_{vec}\frac{R+1}{N}$ (3.12)
with $\tau_{c}$ the typical time scale for evolution of chimeras, $2p_{0}=1$ as the probability of moving in each direction, and where $f(\gamma(t))$ accounts for a normalized factor that decreases the probability of displacement as if it was a random walker in an increasingly crowded environment of positive spins (see Section 3.2.3), and therefore decreasing the probability of one chimera to collide with other chimeras. It is sufficient for the purpose of present analysis to assume that it is a decreasing function of $\gamma(t)$ for $\gamma(t)\in(0,1)$ and satisfies $f(1)=0$, as that would be the attractor scenario. The decrease of the number of chimeras also will depend on the number of pairs of neighboring chimeras at given time, i.e. $N_{c}(t)n_{vec}$, with $n_{vec}$ being the number of neighboring chimeras to a given one (with possible values $n_{nec}=0,1,2$) and on the fraction of the ring that a chimera occupies, i.e. $\frac{R+1}{N}.$
@@
$\tau_{a}\frac{dN_{a}(t)}{dt}=p_{0}f(\gamma(t))N_{c}n_{vec}\frac{R+1}{N}$ (3.13)
with $\tau_{a}$ the time constant for the evolution of attractors. We can see that $N_{a}(t)$ will affect indirectly the dynamics of chimeras through $\gamma(t)$ if we assume that for any time $N_{c}(0)\gamma_{0}=N_{a}(t)+N_{c}(t)\gamma(t)=\bar{\gamma}N$ (see above).
Assuming the choice $f(\gamma(t))=1-\gamma(t),$ Eqs. 3.12 and 3.13 become
@@
$\begin{array}[]{ll}\displaystyle N_{c}(t)=\frac{2\tau_{a}S(t)-\tau_{c}N_{0}}{2\tau_{a}-\tau_{c}}\\ \\ \displaystyle N_{a}(t)=\frac{\tau_{c}\left[N_{0}-S(t)\right]}{2\tau_{a}-\tau_{c}}\end{array}$ (3.15)
@@
$\begin{array}[]{lll}\displaystyle N_{c}(t)&=&\displaystyle N_{0}\,e^{\alpha t}\\ \\ \displaystyle N_{a}(t)&=&\displaystyle\frac{\tau_{c}N_{0}}{2\tau_{a}}\left(1-e^{\alpha t}\right).\end{array}$ (3.16)
@@
Figure 5: Main plot: Dependence of the mean number of chimeras $N_{c}$ surviving at observation time $t_{obs}$ as a function of the system size NN. Each data point has been obtained after averaging the measured $N_{c}$ over $10$ different simulations. Error bars have been computed using the standard deviation of these data. The figure shows a clear scaling as $N_{c}\approx\frac{q_{0}}{\gamma(t_{obs})(R+1)}N$ confirming the theoretical prediction in Eq. 3.9 assuming $\gamma(t_{obs})=0.5$. Parameters $t_{obs}=1000$ MCS, $q_{0}=0.1,$ $R=100$ and system sizes $N=500,1000,2000,5000,10000,20000$ and $40000$ spins. Inset: Slow decrease of $N_{c}$ when the observation time $t_{obs}$ increases from 11 to $2\times 10^{4}$ MCS for a system size of $N=10^{4}$. We observe a good agreement between simulation data and theoretical prediction in Eqs. 3.15 for $\tau_{c}=100$ and $\tau_{a}=200$ MCS. Each data point has been obtained after averaging over $10$ simulations.

### v2

@@ Abstract
[…]
[1]Institute “Carlos I” for Theoretical and Computational Physics, and Department of Electromagnetism and Physics of the Matter, University of Granada, E-18071 Granada, Spain
@@ 1 Introduction
Our study shows that a simple Ising chain with non-local diffusion exhibits a rich phenomenology, including chimera states. In this scenario, we also show that these states constitute the equilibrium state satisfying detailed balance, for certain regions of the parameter space. The present work could be extended to some non-equilibrium situations, including, e.g., competing reaction and diffusion dynamics, a scenario more commonly observed in many social, ecological and biological systems. In this last case, a prominent example is the activity of the mammalian brain, where neural excitability is driven by chemical synapses (reaction) and voltage propagation occurs via gap junctions or electrical synapses among neurons (diffusion) [21, 11, 27]. Then our model can be easily extended to describe brain activity at multiple scales. At the microscopic level, spin variables represent neuronal voltage states, firing or not, with diffusion linked to the presence of electrical synapses in the neural tissue. At mesoscopic or macroscopic scales, it could be used to study local field potential diffusion between neighboring brain regions [11, 10].
@@ 2.1 Non-local diffusion dynamics
$\mathcal{H}(\boldsymbol{\sigma})=-\frac{1}{2}\sum_{i,j}J_{ij}\sigma_{i}\sigma_{j}$ (1)
where $J_{ij}=J\times[0<|i-j|\leq R]$ (with $[P]$ indicating the Iverson bracket) indicates constant ferromagnetic interaction strength among spins ($J>0$) within the diffusion range $R$. Note that ferromagnetic interaction strength $J$ among spins is given in units of $k_{B}T.$ An illustration of the system is depicted in Fig. 1.
@@
Due to the presence of the thermal bath, the system evolves according to stochastic dynamics governed by a master equation – with transition matrix $W(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma})$ – which gives the evolution in a characteristic time $\tau$ of the probability $P(\boldsymbol{\sigma},t)$ of finding the spin system in state $\boldsymbol{\sigma}$ at time $t$ [17, 47, 15], and that can be expressed as
$\tau\frac{dP(\boldsymbol{\sigma},t)}{dt}=\sum\limits_{\boldsymbol{\sigma}^{\prime}}[P(\boldsymbol{\sigma}^{\prime},t)W(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma})-P(\boldsymbol{\sigma},t)W(\boldsymbol{\sigma},\boldsymbol{\sigma}^{\prime})]$ (2)
$W(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma})=\sum\limits_{\begin{subarray}{c}i,j\\ 0<|i-j|\leq R\end{subarray}}\delta_{\sigma_{1},\sigma^{\prime}_{1}}...\delta_{\sigma_{i}\sigma^{\prime}_{j}}...\delta_{\sigma_{j}\sigma^{\prime}_{i}}...\delta_{\sigma_{N},\sigma^{\prime}_{N}}w^{ij}(\boldsymbol{\sigma})$ (3)
@@
$w^{ij}(\boldsymbol{\sigma})=\begin{cases}1\quad\mathrm{if}\;\Delta\mathcal{H}^{ij}\leq 0\\ e^{-\beta\Delta\mathcal{H}^{ij}}\quad\mathrm{if}\;\Delta\mathcal{H}^{ij}>0\ \end{cases}\quad\quad\Delta\mathcal{H}^{ij}=\mathcal{H}(\boldsymbol{\sigma}^{ij})-\mathcal{H}(\boldsymbol{\sigma})$ (4)
@@ 2.2 Simulations and numerical details
We now describe the Monte Carlo algorithm used to simulate our system for an extensive combination of model parameters. The Ising model itself has no intrinsic dynamic and it is when we couple the system to a heat bath at temperature $T$ that stochastic changes in the spin variables occurs so we have a stochastic kinetic version of Ising model driven by the master equation as those in Eq. (3). In a spin system of size $N$, each Monte Carlo Step (MCS) consists of $N$ spin-flip or spin exchange attempts. In general, we have performed simulations of $5\times 10^{4}$ MCS. Each simulation of our system operates according to the following algorithm:
@@
    The Metropolis rate $w^{ij}(\boldsymbol{\sigma})$ is then evaluated following the algorithm described in Eq. (4) and we compute $\Delta\mathcal{H}$ explicitly as follows:
    $\Delta\mathcal{H}^{ij}=\left(\sum\limits_{\begin{subarray}{c}\\ k\neq j\end{subarray}}J_{ik}\sigma_{k}-\sum\limits_{\begin{subarray}{c}k^{\prime}\neq i\end{subarray}}J_{jk^{\prime}}\sigma_{k^{\prime}}\right)\left(\sigma_{i}-\sigma_{j}\right)$ (5)
@@
Spin pairs in the ring are updated using a sequential updating, i.e., at each time step during the evolution of the system one chooses a spin pair at random and exchanges their state using the Metropolis algorithm. Then the system state is updated after each attempt. Note that this is different to what occurs during parallel updating in which, for example in the case of spin-flip dynamics, one performs $N$ spin-flip attempts and updates the state of the whole system state only after these $N$ attempts. Although work regarding parallel updating in the Ising model has been conducted [22] and proven to obey equilibrium statistics, one has to be careful with its implementation so that it satisfies detailed balance [31].
@@ 3.1 Analytical results
$\Delta\mathcal{H}^{ij}=\left(\sum\limits_{\begin{subarray}{c}k=i-R\\ k\neq i,j\end{subarray}}^{i+R}\sigma_{k}-\sum\limits_{\begin{subarray}{c}k^{\prime}=j-R\\ k^{\prime}\neq i,j\end{subarray}}^{j+R}\sigma_{k^{\prime}}\right)\left(\sigma_{i}-\sigma_{j}\right)$ (6)
where we have just computed the terms in $\Delta\mathcal{H}$ summing over $R$ neighbors on each side and considered $J=1$ for simplicity. Note that exchange energy $\Delta\mathcal{H}^{ij}$ is only non-zero if the exchanged spins $i$ and $j$ have different sign.
@@ 3.1.1 Emergence and stability of chimera states
$B_{n^{+}}(i_{1},\ldots,i_{n^{+}})=\left\{k\in\left\{1,\ldots,N\right\},\;k\in\bigcap_{m=1}^{n^{+}}B_{1}(i_{m}),\;|i_{1}-i_{n+}|\leq R\right\}$ (7)
@@
$p_{n^{+}}^{exchange}=\frac{2n^{+}}{N^{2}}|B_{n^{+}}(i_{1},...\,,i_{n^{+}})|\;\;\;\;\;\;\;\;\;\;\;\;p_{n^{+}}^{c}=1-p_{n^{+}}^{exchange}.$ (8)
@@
$W\bigl(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma}\bigr)\;=\;\frac{p_{n^{+}}^{exchange}}{n^{+}}\sum\limits_{k=1}^{n^{+}}\left[\sum\limits_{i_{k},j\in B_{n^{+}}(i_{1},...,i_{n^{+}})}Ex_{i_{k},j}(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma}\bigr)\right]+p_{n^{+}}^{c}\;\Delta(\boldsymbol{\sigma}^{\prime},\boldsymbol{\sigma}\bigr)$ (9)
@@
$\varSigma^{n^{+}}:=\left\{\boldsymbol{\sigma}\in\Lambda\;/\;|i_{1}-i_{n^{+}}|_{p}\leq R,\;i_{1}<i_{2}<\ldots<i_{n^{+}}\right\}.$ (10)
@@
Also, note that the actual construction of $B_{n^{+}}(i_{1},...\,,i_{n^{+}})$ is not always possible, as can be see from definition (7). This follows from the fact that the interaction range $R$ limits the maximum number of positive spins that can simultaneously located within the range of another one. In fact, for configurations with $n^{+}>R+1\Rightarrow|i_{1}-i_{n^{+}}|_{p}>R\Rightarrow\;B_{n^{+}}=\emptyset$. In this situation, there is no rule to determine which spins will have zero energy transitions available. Actually, even for a configuration with $n^{+}=R+1$, regardless of whether $|i_{1}-i_{n^{+}}|_{p}\leq R$, the ball would still be the empty set. This is because any individual ball $B(i_{k})$ (from which we will construct $B_{n^{+}}$) is defined without the index $i_{k}$. This does not imply that exchanges are impossible, but that they are trivial as all spins are positive. We will try to discuss briefly what happens for $n^{+}>R+1$ in the next section.
@@ 3.1.2 Disappearance of chimera states and formation of attractors
Nonetheless, we have arrived to the energy minimum ($\boldsymbol{\sigma}\in\Sigma^{R}$), though any possible exchange would not change the configuration. Following equation (9):
@@
Let us now extend the number of positive spins considering now $n^{+}>R+1$. In this situation, $B_{n^{+}}=\emptyset$ as in the last case, but an important difference is that configurations in the energy minima will not be a configuration of type $\boldsymbol{\sigma}\in\Sigma^{n^{+}}$, since for $n^{+}>R+1$ one has $|i_{1}-i_{n^{+}}|_{p}>R$, so no configuration satisfies the condition to be in $\Sigma^{n^{+}}$, from definition (10). The system will anyway advance stochastically to an energy minimum, but this minimum will not be in any class of type $\Sigma^{X}$. However, we can give an argument of how these minima behave.
@@
To deduct the formation of these attractors, we have supposed that we have a configuration of type $\Pi$ from the start (see proposition 3.1.1). This will obviously not always be the case, and to characterize any kind of general behavior we have to employ a more general approach. In fact, we realized that the Hamiltonian (1) is a special case of a one‐dimensional Ising model with Kac‐type interactions [7, 6, 38]. For these, it is shown that regions of constant magnetization $m$ and m-m alternate throughout the system (what we have called attractors), and that they are of size $L_{attractor}\approx e^{cR}$ for some suitable constant $c$, which we will numerically calculate in the next section.
@@ 3.2 Numerical results for T=0
$m_{0}>\frac{2(R+1)}{N}-1=2r-1+{\cal{O}}\left(N^{-1}\right)=m_{c}$ (12)
Then, if we see that chimeras and attractors coexist in such conditions, this implies that that chimera state will be metastable, as chimeras will always be able to merge into attractors when (12) holds (see also next section).
@@ 3.2.1 Length of attractors
Figure 3 plots $\ell_{dom}$ and $\tau_{ss}$ against $r$ for $m_{0}=0.4$. The figure illustrates that there are multiple $\ell_{dom}$ points for each value of $r$, which correspond to different-seeded runs. The scatter of low $\ell_{dom}$ points across nearly all $r$ values in Fig. 3 signals the intermittent presence of small, noise-driven domains, precisely the hallmark of chimera–attractor coexistence seen in Fig. 2. We fitted $\ell_{dom}$ versus $r$ to two candidate curves:
@@
Preliminary results in Fig. 4 show that the number of chimeras increase with $N$ and that through collisions they can form other chimeras or attractors. This last case occurs due to the fulfillment of condition (12) when $n^{+}$ is increased. So an interesting question to address is how the number of initials chimeras evolve in time until only attractors remain. We have seen that attractors start to emerge when $n^{+}\approx R+1.$ Note moreover that $\langle n^{+}\rangle\approx(1+m_{0})N/2\equiv q_{0}N$, Then, the condition for an attractor to emerge approximately is $q_{0}\approx(R+1)/N.$
@@
$N_{0}\approx\frac{q_{0}}{\gamma_{0}(R+1)}N\equiv\frac{\bar{\gamma}}{\gamma_{0}}N,$ (14)
@@
$L_{chimera}\approx R+1$ (15)
independently of $\gamma_{0}.$ This is consistent with definition in Eq. (11) where now and in general for any time one has $|B_{n^{+}_{c}}(i_{1},\ldots,i_{n^{+}_{c}})|=(1-\gamma(t))(R+1)$ and $n^{+}_{c}=\gamma(t)(R+1)$ We can conclude that initially
$N_{c}(0)=N_{0}\approx\frac{q_{0}}{\gamma_{0}L_{chimera}}N.$ (16)
@@
Under these assumptions, we have tested the Eq. (14) numerically in Fig. 5 (Main plot) for a relatively short observational time of $t_{obs}=1000$ MCS. The figure shows a clear scaling of the number of surviving chimeras at this time with the system size, while its inset indicates that when the observation time increases the number of existing chimeras decreases. Note, moreover, that the initial number of chimeras in the thermodynamic limit can already be established by Eq. (14) just taking $R=rN$ which yields $N_{0}\approx q_{0}/\gamma_{0}r$, expression that only depends on intensive quantities.
@@
$\tau_{c}\frac{dN_{c}(t)}{dt}=-2p_{0}f(\gamma(t))N_{c}(t)n_{vec}\frac{R+1}{N}$ (17)
with $\tau_{c}$ the typical time scale for evolution of chimeras. This equation gives the decrease of the number of chimeras in time (see the negative sign in the right-hand side of the equation). Since a decrease in the number of chimeras mainly occurs because collisions between chimeras generate denser chimeras or attractors, it will depend first in how fast chimeras move in both directions of the ring (this is why the factor 2 appears after the minus) which is related with the probability $p_{0}=1/2$ of moving in each direction of the ring. This probability is multiplied by the factor $f(\gamma(t))$ which accounts for a normalized factor that decreases the probability of displacement $p_{0}$ as if it was a random walker in an increasingly crowded environment of positive spins (see Section 3.2.3), and therefore decreasing the probability of one chimera to collide with other chimeras. It is sufficient for the purpose of present analysis to assume that it is a decreasing function of $\gamma(t)$ for $\gamma(t)\in(0,1)$ and satisfies $f(1)=0$, as that would be the attractor scenario. Finally, the decrease of the number of chimeras also will depend on the number of pairs of neighboring chimeras at given time, i.e. the factor $N_{c}(t)n_{vec}$ in the right-hand side of equation (17), with $n_{vec}$ being the number of neighboring chimeras to a given one (with possible values $n_{nec}=0,1,2$) and on the fraction of the ring that a chimera occupies, i.e. the last factor $\frac{R+1}{N}$ on the right-hand side of Eq. (17).
@@
$\tau_{a}\frac{dN_{a}(t)}{dt}=p_{0}f(\gamma(t))N_{c}n_{vec}\frac{R+1}{N}$ (18)
with $\tau_{a}$ the time constant for the evolution of attractors. The reason for the choice in the right-hand of Eq. (18) is that the number of atractors increase (possitive right hand side of the equation) in the same way that chimeras decreases but its number increases at a rate that is just half of the decrease of chimeras since an attractor generally originates when two chimeras collide. One has to define also a time scale of the evolution of attractor $\tau_{a}$ different than the time scale for the evolution of the chimeras $\tau_{c}$ since not all chimera collisions result in an attractor so $\tau_{a}>\tau_{c}$. We can see that $N_{a}(t)$ will affect indirectly the dynamics of chimeras through $\gamma(t)$ if we assume that for any time $N_{c}(0)\gamma_{0}=N_{a}(t)+N_{c}(t)\gamma(t)=\bar{\gamma}N$ (see above).
Assuming the choice $f(\gamma(t))=1-\gamma(t),$ Eqs. (17) and (18) become
@@
$\begin{array}[]{ll}\displaystyle N_{c}(t)=\frac{2\tau_{a}S(t)-\tau_{c}N_{0}}{2\tau_{a}-\tau_{c}}\\ \\ \displaystyle N_{a}(t)=\frac{\tau_{c}\left[N_{0}-S(t)\right]}{2\tau_{a}-\tau_{c}}\end{array}$ (20)
@@
$\begin{array}[]{lll}\displaystyle N_{c}(t)&=&\displaystyle N_{0}\,e^{\alpha t}\\ \\ \displaystyle N_{a}(t)&=&\displaystyle\frac{\tau_{c}N_{0}}{2\tau_{a}}\left(1-e^{\alpha t}\right).\end{array}$ (21)
@@
Figure 5: Main plot: Dependence of the mean number of chimeras $N_{c}$ surviving at observation time $t_{obs}$ as a function of the system size NN. Each data point has been obtained after averaging the measured $N_{c}$ over $10$ different simulations. Error bars have been computed using the standard deviation of these data. The figure shows a clear scaling as $N_{c}\approx\frac{q_{0}}{\gamma(t_{obs})(R+1)}N$ confirming the theoretical prediction in Eq. (14) assuming $\gamma(t_{obs})=0.5$. Parameters $t_{obs}=1000$  MCS, $q_{0}=0.1,$ $R=100$ and system sizes $N=500,1000,2000,5000,10000,20000$ and $40000$ spins. Inset: Slow decrease of $N_{c}$ when the observation time $t_{obs}$ increases from 11 to $2\times 10^{4}$ MCS for a system size of $N=10^{4}$. We observe a good agreement between simulation data and theoretical prediction in Eqs. (20) for $\tau_{c}=100$ and $\tau_{a}=200$ MCS. Each data point has been obtained after averaging over $10$ simulations.